OpenStudy (anonymous):

Sorry, the equation's really small /: Simplify the complex rational expression

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x+6 } }{ \frac{ 5 }{ x^2-36 } }\]

OpenStudy (anonymous):

I think I have to put them beside each other

OpenStudy (luigi0210):

Take the bottom and flip it

OpenStudy (anonymous):

\[ \frac{ x^2-36 }{ 5 }\] like that?

OpenStudy (luigi0210):

\[\LARGE \frac{ \frac{ 1 }{ x+6 } }{ \frac{ 5 }{ x^2-36 } }=\frac{1}{x+6}*\frac{x^2-36}{5}\] Now solve it

OpenStudy (anonymous):

\[\frac{ -x^2 +36 }{ 5x+30 }\] is that correct?

OpenStudy (anonymous):

wrong question sorry lol ignore that

OpenStudy (luigi0210):

You actually could have simplified earlier, if you notice, \(x^2-36\) factors to \((x-6)(x+6)\) . Cross out the \((x+6)\) and you are left with \(\Large \frac{x-6}{5}\)

OpenStudy (anonymous):

I always forget when to factor...I've never really been good a that but that seems simpler... than simplifying at the last minute. Thank you.